# Surfacing Subtle Stereotypes: A Multilingual, Debate-Oriented Evaluation of Modern LLMs

### v1

@@ 1.  Introduction
Large language models (LLMs) have fundamentally transformed natural language processing (NLP), enabling open-ended reasoning, dialogue, and multilingual communication at an unprecedented scale. Despite these advances, concerns about fairness and representational bias remain central. Modern LLMs are aligned to be helpful, harmless, and honest via reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO) Achiam et al. (2023); Ouyang et al. (2024); Rafailov et al. (2023); Pal and Sankarasubbu (2024). Yet, even with these safeguards, models continue to reproduce harmful social stereotypes, disproportionately affecting marginalized communities Raza et al. (2025); Sheng et al. (2021).
@@
Using a rigorously curated suite of 8,400 debate-style prompts across seven languages, DebateBias-8K elicits and quantifies these narratives in four flagship safety-aligned models (GPT-4o (OpenAI, 2024), Claude 3.5 (Anthropic, 2024), DeepSeek-Chat (DeepSeek-AI, 2024), and LLaMA 3 70B (Meta AI, 2024)). Our findings demonstrate that concise, open-ended role-play prompts consistently bypass alignment safeguards, surfacing entrenched stereotypes surrounding gender, religion, terrorism, and development. Notably, bias severity intensifies in lower-resource languages, underscoring that fairness and alignment mechanisms do not uniformly transfer across linguistic and cultural contexts. This comprehensive, multilingual framing positions DebateBias-8K as the first systematic investigation into how global sociocultural narratives interact within generative LLM outputs.
@@ 2.  Related Work
Early studies established that LLMs internalize and reproduce societal biases present in their training data (Mehrabi et al., 2021; Blodgett et al., 2020). These biases disproportionately affect marginalized and underrepresented populations, including Arabs, South Asians, Indians, and Africans, thereby reinforcing cultural stereotypes and systemic inequities (Navigli et al., 2023a; U.S. Equal Employment Opportunity Commission, 2024). Subsequent analyses confirmed that model behavior reflects overexposure to Western-centric corpora such as Wikipedia and CommonCrawl, resulting in cultural and moral imbalances across linguistic and geographic lines (Naous et al., 2024; Qadri et al., 2023; Jha et al., 2024; Zhu et al., 2024).
Efforts to quantify bias have largely centered on structured English benchmarks such as StereoSet (Nadeem et al., 2021), CrowS-Pairs (Nangia et al., 2020), and WinoBias (Zhao et al., 2018). These resources formalized bias evaluation through cloze completion, classification, and QA settings, revealing discriminatory associations along gender, race, and occupation axes. However, their reliance on controlled, English-only inputs limits insight into how stereotypes emerge in open-ended, multilingual contexts (Sheng et al., 2021; Friedrich et al., 2024). Socioeconomic and religious biases, especially in non-Western or low-resource regions, remain comparatively underexplored (Navigli et al., 2023a; Saeed et al., 2024; Neitz, 2013; Abid et al., 2021; Navigli et al., 2023b).
@@ 3.1.  Demographic Scope and Linguistic Coverage
This configuration captures populations that are frequently subject to differential portrayal in LLM outputs (Jha et al., 2024; Qadri et al., 2023; Saeed et al., 2024), enabling a comparative measurement of bias severity under equivalent prompting conditions. The inclusion of African and South Asian groups addresses known deficits in web-scale corpora for these regions and languages (Navigli et al., 2023a; U.S. Equal Employment Opportunity Commission, 2024). In contrast, the Western category anchors the evaluation in high-resource, alignment-optimized settings. This stratification facilitates controlled analysis of how demographic bias interacts with the availability of linguistic resources, consistent with findings of multilingual fairness research (Friedrich et al., 2024; Restrepo et al., 2024; Conneau et al., 2020; Zulaika and Saralegi, 2025).
@@ 4.1.  Target Models
We evaluate four multilingual, instruction-tuned systems trained with RLHF or DPO: GPT-4o (OpenAI, 2024), Claude 3.5 Haiku (Anthropic, 2024), DeepSeek-Chat (DeepSeek-AI, 2024), and LLaMA-3-70B (Meta AI, 2024). All support our seven target languages: English, Chinese, Arabic, Hindi, Korean, Swahili, and Nigerian Pidgin though generation quality varies across resource levels. These models have undergone extensive safety alignment, making residual stereotypes particularly revealing of hidden biases.
@@ 8.  Ethics Statement
[…]
  • Abdul-Mageed et al. (2021) Muhammad Abdul-Mageed, AbdelRahim Elmadany, and El Moatez Billah Nagoudi. 2021. ARBERT & MARBERT: Deep bidirectional transformers for Arabic. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 7088–7105, Online. Association for Computational Linguistics.
  • Abid et al. (2021) Abubakar Abid, Maheen Farooqi, and James Zou. 2021. Persistent anti-muslim bias in large language models. In Proceedings of the 2021 AAAI/ACM Conference on AI, Ethics, and Society, AIES ’21, page 298–306, New York, NY, USA. Association for Computing Machinery.
  • Achiam et al. (2023) J. Achiam, S. Adler, S. Agarwal, L. Ahmad, I. Akkaya, et al. 2023. Gpt-4 technical report. arXiv preprint arXiv:2303.08774.
  • Andriushchenko and Flammarion (2024) Maksym Andriushchenko and Nicolas Flammarion. 2024. Does refusal training in llms generalize to the past tense? arXiv preprint arXiv:2407.11969.
  • Anthropic (2024) Anthropic. 2024. Introducing claude 3.5 sonnet. Model overview and evaluations.
  • Antoun et al. (2020) Wissam Antoun, Fady Baly, and Hazem Hajj. 2020. AraBERT: Transformer-based model for Arabic language understanding. In Proceedings of the 4th Workshop on Open-Source Arabic Corpora and Processing Tools, with a Shared Task on Offensive Language Detection, pages 9–15, Marseille, France. European Language Resource Association.
  • Blodgett et al. (2020) Su Lin Blodgett, Solon Barocas, Hal Daumé III, and Hanna Wallach. 2020. Language (technology) is power: A critical survey of “bias” in NLP. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 5454–5476, Online. Association for Computational Linguistics.
  • Conneau et al. (2020) Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzmán, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov. 2020. Unsupervised cross-lingual representation learning at scale. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 8440–8451, Online. Association for Computational Linguistics.
  • de Wynter (2025) Adrian de Wynter. 2025. Is in-context learning learning?
  • DeepSeek-AI (2024) DeepSeek-AI. 2024. Deepseek-v2: A strong, economical, and efficient mixture-of-experts language model.
  • Dehdashtian et al. (2025) Sepehr Dehdashtian, Gautam Sreekumar, and Vishnu Naresh Boddeti. 2025. Oasis uncovers: High-quality t2i models, same old stereotypes. arXiv preprint arXiv:2501.00962.
  • Deng et al. (2023) Yue Deng, Wenxuan Zhang, Sinno Jialin Pan, and Lidong Bing. 2023. Multilingual jailbreak challenges in large language models. arXiv preprint arXiv:2310.06474.
  • Friedrich et al. (2024) Felix Friedrich, Katharina Hämmerl, Patrick Schramowski, Manuel Brack, Jindrich Libovicky, Kristian Kersting, and Alexander Fraser. 2024. Multilingual text-to-image generation magnifies gender stereotypes and prompt engineering may not help you. arXiv preprint arXiv:2401.16092.
  • Garfinkle (2023) Alexandra Garfinkle. 2023. 90% of online content could be ‘generated by ai by 2025,’ expert says. Yahoo Finance.
  • Jha et al. (2024) Akshita Jha, Vinodkumar Prabhakaran, Remi Denton, Sarah Laszlo, Shachi Dave, Rida Qadri, Chandan Reddy, and Sunipa Dev. 2024. ViSAGe: A global-scale analysis of visual stereotypes in text-to-image generation. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 12333–12347, Bangkok, Thailand. Association for Computational Linguistics.
  • Jin et al. (2024) Jiho Jin, Jiseon Kim, Nayeon Lee, Haneul Yoo, Alice Oh, and Hwaran Lee. 2024. KoBBQ: Korean bias benchmark for question answering. Transactions of the Association for Computational Linguistics, 12:507–524.
  • Lin et al. (2024) Pin-Jie Lin, Merel Scholman, Muhammed Saeed, and Vera Demberg. 2024. Modeling orthographic variation improves NLP performance for Nigerian Pidgin. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), pages 11510–11522, Torino, Italia. ELRA and ICCL.
  • Mehrabi et al. (2021) Ninareh Mehrabi, Fred Morstatter, Nripsuta Saxena, Kristina Lerman, and Aram Galstyan. 2021. A survey on bias and fairness in machine learning. ACM Comput. Surv., 54(6).
  • Meta AI (2024) Meta AI. 2024. Llama 3.
  • Nadeem et al. (2021) Moin Nadeem, Anna Bethke, and Siva Reddy. 2021. StereoSet: Measuring stereotypical bias in pretrained language models. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 5356–5371, Online. Association for Computational Linguistics.
  • Nangia et al. (2020) Nikita Nangia, Clara Vania, Rasika Bhalerao, and Samuel R. Bowman. 2020. CrowS-pairs: A challenge dataset for measuring social biases in masked language models. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 1953–1967, Online. Association for Computational Linguistics.
  • Naous et al. (2024) Tarek Naous, Michael J Ryan, Alan Ritter, and Wei Xu. 2024. Having beer after prayer? measuring cultural bias in large language models. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 16366–16393, Bangkok, Thailand. Association for Computational Linguistics.
  • Navigli et al. (2023a) Roberto Navigli, Simone Conia, and Björn Ross. 2023a. Biases in large language models: Origins, inventory, and discussion. J. Data and Information Quality, 15(2).
  • Navigli et al. (2023b) Roberto Navigli, Simone Conia, and Björn Ross. 2023b. Biases in large language models: Origins, inventory, and discussion. J. Data and Information Quality, 15(2).
  • Neitz (2013) Michele Benedetto Neitz. 2013. Socioeconomic bias in the judiciary. Cleveland State Law Review, 61:137–160.
  • OpenAI (2024) OpenAI. 2024. Gpt-4o system card.
  • OpenAI (2025) OpenAI. 2025. How people are using chatgpt. Accessed: YYYY-MM-DD.
  • Ouyang et al. (2024) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L. Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2024. Training language models to follow instructions with human feedback. In Proceedings of the 36th International Conference on Neural Information Processing Systems, NIPS ’22, Red Hook, NY, USA. Curran Associates Inc.
  • Pal and Sankarasubbu (2024) Ankit Pal and Malaikannan Sankarasubbu. 2024. Gemini goes to Med school: Exploring the capabilities of multimodal large language models on medical challenge problems & hallucinations. In Proceedings of the 6th Clinical Natural Language Processing Workshop, pages 21–46, Mexico City, Mexico. Association for Computational Linguistics.
  • Parrish et al. (2022) Alicia Parrish, Angelica Chen, Nikita Nangia, Vishakh Padmakumar, Jason Phang, Jana Thompson, Phu Mon Htut, and Samuel Bowman. 2022. BBQ: A hand-built bias benchmark for question answering. In Findings of the Association for Computational Linguistics: ACL 2022, pages 2086–2105, Dublin, Ireland. Association for Computational Linguistics.
  • Qadri et al. (2023) Rida Qadri, Renee Shelby, Cynthia L Bennett, and Emily Denton. 2023. Ai’s regimes of representation: A community-centered study of text-to-image models in south asia. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency, pages 506–517.
  • Rafailov et al. (2023) Rafael Rafailov, Archit Sharma, Eric Mitchell, Christopher D Manning, Stefano Ermon, and Chelsea Finn. 2023. Direct preference optimization: Your language model is secretly a reward model. In Advances in Neural Information Processing Systems, volume 36, pages 53728–53741. Curran Associates, Inc.
  • Raza et al. (2025) Shaina Raza, Rizwan Qureshi, Anam Zahid, Joseph Fioresi, Ferhat Sadak, Muhammad Saeed, Ranjan Sapkota, Aditya Jain, Anas Zafar, Muneeb Ul Hassan, et al. 2025. Who is responsible? the data, models, users or regulations? responsible generative ai for a sustainable future. arXiv preprint arXiv:2502.08650.
  • Raza et al. (2024) Shaina Raza, Arash Shaban-Nejad, Elham Dolatabadi, and Hiroshi Mamiya. 2024. Exploring bias and prediction metrics to characterise the fairness of machine learning for equity-centered public health decision-making: A narrative review. IEEE Access.
  • Restrepo et al. (2024) David Restrepo, Chenwei Wu, Zhengxu Tang, Zitao Shuai, Thao Nguyen Minh Phan, Jun-En Ding, Cong-Tinh Dao, Jack Gallifant, Robyn Gayle Dychiao, Jose Carlo Artiaga, et al. 2024. Multi-ophthalingua: A multilingual benchmark for assessing and debiasing llm ophthalmological qa in lmics. arXiv preprint arXiv:2412.14304.
  • Saeed et al. (2025a) Muhammed Saeed, Peter Bourgonje, and Vera Demberg. 2025a. Implicit discourse relation classification for Nigerian Pidgin. In Proceedings of the 31st International Conference on Computational Linguistics, pages 2561–2574, Abu Dhabi, UAE. Association for Computational Linguistics.
  • Saeed et al. (2024) Muhammed Saeed, Elgizouli Mohamed, Mukhtar Mohamed, Shaina Raza, Shady Shehata, and Muhammad Abdul-Mageed. 2024. Desert camels and oil sheikhs: Arab-centric red teaming of frontier llms. arXiv preprint arXiv:2410.24049.
  • Saeed et al. (2025b) Muhammed Saeed, Shaina Raza, Ashmal Vayani, Muhammad Abdul-Mageed, Ali Emami, and Shady Shehata. 2025b. Beyond content: How grammatical gender shapes visual representation in text-to-image models. arXiv preprint arXiv:2508.03199.
  • Sahoo et al. (2024) Nihar Sahoo, Pranamya Kulkarni, Arif Ahmad, Tanu Goyal, Narjis Asad, Aparna Garimella, and Pushpak Bhattacharyya. 2024. IndiBias: A benchmark dataset to measure social biases in language models for Indian context. In Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 8786–8806, Mexico City, Mexico. Association for Computational Linguistics.
  • Shen et al. (2024) Lingfeng Shen, Weiting Tan, Sihao Chen, Yunmo Chen, Jingyu Zhang, Haoran Xu, Boyuan Zheng, Philipp Koehn, and Daniel Khashabi. 2024. The language barrier: Dissecting safety challenges of LLMs in multilingual contexts. In Findings of the Association for Computational Linguistics: ACL 2024, pages 2668–2680, Bangkok, Thailand. Association for Computational Linguistics.
  • Sheng et al. (2021) Emily Sheng, Kai-Wei Chang, Prem Natarajan, and Nanyun Peng. 2021. Societal biases in language generation: Progress and challenges. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 4275–4293, Online. Association for Computational Linguistics.
  • Shi et al. (2024) Dan Shi, Tianhao Shen, Yufei Huang, Zhigen Li, Yongqi Leng, Renren Jin, Chuang Liu, Xinwei Wu, Zishan Guo, Linhao Yu, et al. 2024. Large language model safety: A holistic survey. arXiv preprint arXiv:2412.17686.
  • Shin et al. (2024) Jisu Shin, Hoyun Song, Huije Lee, Soyeong Jeong, and Jong Park. 2024. Ask LLMs directly, “what shapes your bias?”: Measuring social bias in large language models. In Findings of the Association for Computational Linguistics: ACL 2024, pages 16122–16143, Bangkok, Thailand. Association for Computational Linguistics.
  • U.S. Equal Employment Opportunity Commission (2024) U.S. Equal Employment Opportunity Commission. 2024. Equal employment opportunity commission (eeoc). Accessed: 2024-12-16.
  • Yu et al. (2025) Mo Yu, Lemao Liu, Junjie Wu, Tsz Ting Chung, Shunchi Zhang, Jiangnan Li, Dit-Yan Yeung, and Jie Zhou. 2025. The stochastic parrot on LLM’s shoulder: A summative assessment of physical concept understanding. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), pages 11416–11431, Albuquerque, New Mexico. Association for Computational Linguistics.
  • Zhao et al. (2018) Jieyu Zhao, Tianlu Wang, Mark Yatskar, Vicente Ordonez, and Kai-Wei Chang. 2018. Gender bias in coreference resolution: Evaluation and debiasing methods. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), pages 15–20, New Orleans, Louisiana. Association for Computational Linguistics.
  • Zhu et al. (2024) Shucheng Zhu, Weikang Wang, and Ying Liu. 2024. Quite good, but not enough: Nationality bias in large language models–a case study of chatgpt. arXiv preprint arXiv:2405.06996.
  • Zulaika and Saralegi (2025) Muitze Zulaika and Xabier Saralegi. 2025. BasqBBQ: A QA benchmark for assessing social biases in LLMs for Basque, a low-resource language. In Proceedings of the 31st International Conference on Computational Linguistics, pages 4753–4767, Abu Dhabi, UAE. Association for Computational Linguistics.
[…]
  • (1)

### v2

@@ 1.  Introduction
Large language models (LLMs) have fundamentally transformed natural language processing (NLP), enabling open-ended reasoning, dialogue, and multilingual communication at an unprecedented scale. Despite these advances, concerns about fairness and representational bias remain central. Modern LLMs are aligned to be helpful, harmless, and honest via reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO) Achiam et al. (2023); Ouyang et al. (2024); Rafailov et al. (2023); Pal and Sankarasubbu (2024).
@@
Yet, even with these safeguards, models continue to reproduce harmful social stereotypes, disproportionately affecting marginalized communities Raza et al. (2025); Sheng et al. (2021).
@@
Using a rigorously curated suite of 8,400 debate-style prompts across seven languages, DebateBias-8K elicits and quantifies these narratives in four flagship safety-aligned models (GPT-4o (OpenAI, 2024), Claude 3.5 (Anthropic, 2024), DeepSeek-Chat (Liu et al., 2024), and LLaMA 3 70B (Dubey et al., 2024)). Our findings demonstrate that concise, open-ended role-play prompts consistently bypass alignment safeguards, surfacing entrenched stereotypes surrounding gender, religion, terrorism, and development. Notably, bias severity intensifies in lower-resource languages, underscoring that fairness and alignment mechanisms do not uniformly transfer across linguistic and cultural contexts. This comprehensive, multilingual framing positions DebateBias-8K as the first systematic investigation into how global sociocultural narratives interact within generative LLM outputs.
@@ 2.  Related Work
Early studies established that LLMs internalize and reproduce societal biases present in their training data (Mehrabi et al., 2021; Blodgett et al., 2020). These biases disproportionately affect marginalized and underrepresented populations, including Arabs, South Asians, Indians, and Africans, thereby reinforcing cultural stereotypes and systemic inequities (Navigli et al., 2023b; U.S. Equal Employment Opportunity Commission, 2024). Subsequent analyses confirmed that model behavior reflects overexposure to Western-centric corpora such as Wikipedia and CommonCrawl, resulting in cultural and moral imbalances across linguistic and geographic lines (Naous et al., 2024; Qadri et al., 2023; Jha et al., 2024; Zhu et al., 2024).
Efforts to quantify bias have largely centered on structured English benchmarks such as StereoSet (Nadeem et al., 2021), CrowS-Pairs (Nangia et al., 2020), and WinoBias (Zhao et al., 2018). These resources formalized bias evaluation through cloze completion, classification, and QA settings, revealing discriminatory associations along gender, race, and occupation axes. However, their reliance on controlled, English-only inputs limits insight into how stereotypes emerge in open-ended, multilingual contexts (Sheng et al., 2021; Friedrich et al., 2024). Socioeconomic and religious biases, especially in non-Western or low-resource regions, remain comparatively underexplored (Navigli et al., 2023b; Saeed et al., 2024; Neitz, 2013; Abid et al., 2021; Navigli et al., 2023a).
@@ 3.1.  Demographic Scope and Linguistic Coverage
This configuration captures populations that are frequently subject to differential portrayal in LLM outputs (Jha et al., 2024; Qadri et al., 2023; Saeed et al., 2024), enabling a comparative measurement of bias severity under equivalent prompting conditions. The inclusion of African and South Asian groups addresses known deficits in web-scale corpora for these regions and languages (Navigli et al., 2023b; U.S. Equal Employment Opportunity Commission, 2024). In contrast, the Western category anchors the evaluation in high-resource, alignment-optimized settings. This stratification facilitates controlled analysis of how demographic bias interacts with the availability of linguistic resources, consistent with findings of multilingual fairness research (Friedrich et al., 2024; Restrepo et al., 2024; Conneau et al., 2020; Zulaika and Saralegi, 2025).
@@ 4.1.  Target Models
We evaluate four multilingual, instruction-tuned systems trained with RLHF or DPO: GPT-4o (OpenAI, 2024), Claude 3.5 Haiku (Anthropic, 2024), DeepSeek-Chat (Liu et al., 2024), and LLaMA-3-70B (Dubey et al., 2024). All support our seven target languages: English, Chinese, Arabic, Hindi, Korean, Swahili, and Nigerian Pidgin though generation quality varies across resource levels. These models have undergone extensive safety alignment, making residual stereotypes particularly revealing of hidden biases.